The product of two positive consecutive integers is 24 more than nine times the larger integer. Find the integers.
Answer by CubeyThePenguin(3113)

consecutive integers: x, (x+1)

x(x+1) = 24 + 9(x+1)
x^2 + x = 9x + 33
x^2 - 8x - 33 = 0
(x - 11)(x + 3) = 0
x = 11, x = -3

x has to be positive, so the integers are 11 and 12.